OpenStudy (anonymous):

How do I use z^n=r^n cis(n(θ+2πk)) to find the exact value of cos(2π/5)?

zepdrix (zepdrix):

hmmmmmmmmm

OpenStudy (anonymous):

So im just researching about de Moivre's Theorem now..

OpenStudy (anonymous):

cis? Does that mean theta was born an angle and is content with being an angle?

zepdrix (zepdrix):

-_-

zepdrix (zepdrix):

I don't even understand what you just said wio :p

OpenStudy (anonymous):

oh cis is just the term we use for (cos theta+isin theta) sorry for the confusion

OpenStudy (anonymous):

Oh ok.

OpenStudy (anonymous):

I suppose you want to let \(r=1\)? Not sure yet.

OpenStudy (anonymous):

\(n=2/5\)?

OpenStudy (anonymous):

im looking at http://answers.yahoo.com/question/index?qid=20081214120934AA11Edn and it looks similar to what im having trouble with

zepdrix (zepdrix):

\[\Large\bf\sf \cos\left(\frac{2\pi}{5}\right)\quad=\quad \operatorname{\Re}\left(\cos\frac{2\pi}{5}+\mathcal i \sin\frac{2\pi}{5}\right)\]Do we want to do this maybe? Use the "real part" notation to get it into complex form. Then maybe we can apply De Moivre's Theorem or something...\[\Large\bf\sf \operatorname{\Re}\left(\cos 2\pi+\mathcal i \sin 2\pi\right)^{1/5}\]Mmmmmmmmmmmmmmmmmm I dunno... Just throwing ideas out.

zepdrix (zepdrix):

I could explain the one in the yahoo answers lol. But I don't quite understand this problem XD Grr

OpenStudy (anonymous):

ah, this website has it all me thinks https://suite101.com/a/de-moivres-theorem-examples-cos3x-sin3x-cos4x-and-sin4x-a271768

OpenStudy (anonymous):

when you expand z^5 It said to use z=x+yi and then expand (x+yi)^5

OpenStudy (anonymous):

Okay, how about this... let \(n = 1/5\).